# Machine Learns Quantum Complexity?

### v1

@@ 1 Introduction
Recently, the concept of quantum complexity [1, 2], which quantifies the difficulty of preparing a given quantum state from a reference state using a specified set of simple (unitary) operations, has received considerable attention [3, 4, 5, 6, 7, 8]. In particular, the Krylov spread complexity, first introduced in [9], is a state-complexity measure that characterizes the spread of a quantum state during its Hamiltonian time evolution. Specifically, it quantifies how the state evolves and spreads within the Krylov subspace, which is spanned by repeated applications of the system’s Hamiltonian on an initial state. (For the Krylov operator complexity, see for instance [10].) In a chaotic system, such as one described by a random Hamiltonian, the Krylov spread complexity exhibits a universal behavior over time [9, 11, 12]. This behavior includes a long ramp, a peak, and a subsequent slope that leads to a plateau. Notably, the ramp extends exponentially long [13, 14], up to a timescale that is exponential in the entanglement entropy of the chaotic system (e.g., $N\sim e^{S}$ in our case below). This universal behavior is closely related to that observed in random matrix theory (RMT) [15], which is further argued to be connected to the late-time dynamics of black holes [16].
@@ 2 Krylov Spread Complexity
In this section, we shall briefly review the basic construct of the Krylov spread complexity, which was first introduced in [9]. Let us start with a quantum system with a Hamiltonian $H$ and the corresponding Hilbert space ${\cal H}$. We are primarily interested in the spread of a time-evolving quantum state
@@
where the weight $n$ may be replaced by any positive increasing sequence of real numbers. In this definition, any other ordered orthonormal basis =${\cal B}=\{|B_{n}\rangle\mid n=0,1,2,\dots\}$ could be used instead of the Krylov basis ${\cal K}$. However it has been shown [9] that the Krylov basis minimizes the complexity (2.4) over all possible choices of such bases $\cal B$. Note that, in this Krylov basis, the Hamiltonian becomes tri-diagonal as
@@
purifies the thermal ensemble where $|n,n\rangle=|n\rangle_{L}\otimes|n\rangle_{R}$ belongs to the left-right doubled Hilbert space ${\cal H}\otimes{\cal H}$ and $Z=\sum_{n}e^{-\beta E_{n}}$ is the canonical partition function. With $H_{R}=1\otimes H$ and $H_{L}=H\otimes 1$, the TFD state is invariant under the boost transformation with a generator $(H_{R}-H_{L})/2$. Hence the $H_{L}$ operation may be equivalently replaced by the $H_{R}$ operation and the general time evolution of the TFD system may be simply given by $e^{-iH_{R}t}$ with $e^{-iH_{R}t}|n,n\rangle=e^{-iE_{n}t}|n,n\rangle$. The left-right system in the TFD state is maximally entangled for a given temperature and dual to the two-sided black hole in the AdS/CFT correspondence [17]. It has been observed [9, 11] that, especially for a chaotic system, there appears a universal behavior of the Krylov spread complexity which exhibits a long ramp, a peak, a subsequent slope leading to a plateau as a function of time. In particular it shows an exponentially long ramp up to the time scale which is exponential in the entanglement entropy of the left-right system for a chaotic Hamiltonian. This universal behavior is closely related to that of the RMT [12] which is also closely related to the late-time behavior of the black hole dynamics [16]. In this note, we shall focus on a finite dimensional RMT with the GUE as a typical example of the Krylov spread complexity of the corresponding TFD state with a chaotic spectrum.
@@ 3.1 Model Architecture
where $o_{i}$ is the $i$-th component of the output in the final layer $L-1$, which has $N_{(L-1)}$ components. Here $l$ is the layer index, ranging from 1 to $L-1$. $h_{j}^{(l)}$ is the value of node $j$ in layer $l$, which has $N_{(l)}$ nodes. The input layer nodes are represented by $h_{n}^{(0)}=V_{n}$, the $n$-th component of the $N_{(0)}$ input variables. $f_{out}$ is the activation function for the output nodes. This function (for example, $Softmax$, $Sigmoid$, or $Linear$) is selected based on the specific problem of interest. Similarly the function $f_{h}$ represents the activation for the nodes $h^{(l)}$ in inner (or hidden) layers $10<l<L-$ 1, where usually a nonlinear function such as $tanh$ or Rectified Linear Unit (ReLU) is chosen for the activation. These nonlinear functions are used to introduce nonlinearity into the model, which is essential for learning complex patterns. Specifically, the function ReLU is defined as $f_{\text{ReLU}}(x)=\alpha x$ if $x>0$, otherwise $f_{\text{ReLU}}(x)=0$, with a typical choice of $\alpha>0$. The connection parameters $W$ and the bias $b$, which are randomly initialized in an appropriate manner [20], represent the weights applied to the output vector of the nodes in the previous layer and the corresponding offset. These parameters will be repeatedly updated and optimized during training iteration by minimizing the loss associated with the task at hand.
@@
where $N_{ch}^{(l-1)}$ is the number of channels (kernels) of layer $l-1$, $n_{1}\,(=1,\cdots,X_{(l)})$ and $n_{2}\,(=1,\cdots,Y_{(l)})$ represent the 2D pixel coordinates on the feature maps with sizes of ($X_{(l)},Y_{(l)}$) in layer $l$. In the case of the input layer, $X_{(0)}$ and $Y_{(0)}$ correspond to the number of pixels along $x$ and $y$ axis of the input image. Assuming no padding is used [^1] Padding refers to the process of adding extra pixels, typically zeros, around the borders of an image (or feature map) to control the spatial size reduction caused by convolution operations. In ‘valid’ padding (i.e. no padding), the size of the feature map shrinks, while ‘same’ padding maintains the input size by adding enough pixels to the borders., the feature map size at layer $l$ follows from the relations $X_{(l)}=X_{(l-1)}-K+1$ and $Y_{(l)}=Y_{(l-1)}-K+1$. $m_{1}$ and $m_{2}$ representing convolution indices take integer values ranging from 00 to $K-1$ with the kernel size $K$. Finally, the indices $i$ and $j$ represent the channel index of layer $l$ and $l-1$ respectively.
@@ 3.2 Training
Our model is trained using a supervised learning approach, where the algorithm learns to map inputs to their corresponding outputs based on labeled training data. Specifically, we employ a dataset of $N\,(=1024)$-dimensional state vectors paired with corresponding quantum complexity values. In this study, the entire dataset consists of $M=200$ samples of GUE, where each sample gives $30723N=$ 3072 state vector data points with $3N$ denoting the total duration of our integer-valued time variable. Among them, 160 samples ( or $160\times 3N$ data points ) are used for training, while $20\times 3N$ data points are used for validation, and another $20\times 3N$ data points are used for testing. The loss function $\mathcal{L}$ used is mean squared error (MSE),
@@
where $C_{I}$ is the predicted quantum complexity value for the $I$-th input data, $\hat{C}_{I}$ is the true quantum complexity value, and batch size $\mathcal{N}$, the number of training examples used in one iteration of the model parameters ($W,b$) update, is set to 32222The batch size 32 (taken as some power of two) is a common choice because it provides a balance between computational efficiency and model convergence. A smaller batch size (like 1) leads to noisy updates and slower convergence, while a larger batch size (equal to the entire dataset) makes the training process inefficient by updating too infrequently. The optimal batch size is usually determined through experimentation, balancing speed and accuracy. . However, when calculating the validation loss, instead of $\mathcal{N}=32$, we used the total number of data points in the validation dataset as the channel number i.e. $\mathcal{N}=20\times 3N=61,440$. Similarly, when calculating the test error, which is the root mean squared error (RMSE), we also use the total number of data points in the test dataset, so $\mathcal{N}=20\times 3N=61,440$. The loss function is minimized with respect to the model parameters $W$ and $b$ using the Adam optimizer [22], an adaptive learning rate optimization algorithm known for its efficiency and effectiveness in training DL models. The model is trained for a total of 100 epochs, where an epoch is one complete pass through the entire training dataset, since the training set has a total of $520160\times 3N=491,$ 520 data points, and the batch size is 32, each epoch corresponds to 15,360 training iterations. In Figure 4, we plot loss functions against epochs for several different bases to examine the convergence behavior of the models during training. The convergence test is an important task in numerical calculations to illustrate the capability of our model and to detect potential problems during training. As can be seen in Figure 4, the stable convergence of loss functions with respect to epochs was obtained for the energy and Krylov bases. Perhaps due to the random nature of the original basis, our training algorithms get stuck immediately. For the pseudo-random basis, both training and validation processes still converge to a much larger value of the loss function compared to those of the energy and Krylov bases. Contrary to expectations, our results suggest that the choice of basis appears to be crucial in applying DL to the complexity.
@@ 5 Conclusions
In this note, we study how a machine based on DL algorithms learns the Krylov spread complexity from patterns of quantum states in systems with $N\times N$ random Hamiltonians drawn from the GUE of matrices. Taking the initial state as the TFD state, we found that the DL algorithm performs well for states at any time, including the plateau region where the states become almost featureless and random.
@@
Our investigation of the Krylov spread complexity in this note may be extended to other systems, such as random matrix theories with other ensembles, field theories [23, 24, 25], non-chaotic systems [26], and so on. Here, we use $N=1024$, but studying the learning efficiency as $N$ increases may also be of interest. However, this would involve computational challenges and warrants further investigation.
@@ Acknowledgement
DB was supported in part by NRF Grant RS-2023-00208011, and by Basic Science Research Program through NRF Grant 2018R1A6A1A06024977 funded by the Ministry of Education. This work was also supported by the UBAI computing resources at the University of Seoul.

### v3

@@ 1 Introduction
Recently, the concept of quantum complexity [1, 2], which quantifies the difficulty of preparing a given quantum state from a reference state using a specified set of simple (unitary) operations, has received considerable attention [3, 4, 5, 6, 7, 8, 9]. In particular, the Krylov spread complexity, first introduced in [10], is a state-complexity measure that characterizes the spread of a quantum state during its Hamiltonian time evolution [11]. Specifically, it quantifies how the state evolves and spreads within the Krylov subspace, which is spanned by repeated applications of the system’s Hamiltonian on an initial state. (For the Krylov operator complexity, see for instance [12].) In a chaotic system, such as one described by a random Hamiltonian, the Krylov spread complexity exhibits a universal behavior over time [10, 13, 14, 15]. This behavior includes a long ramp, a peak, and a subsequent slope that leads to a plateau. Notably, the ramp extends exponentially long [16, 17], up to a timescale that is exponential in the entanglement entropy of the chaotic system (e.g., $N\sim e^{S}$ in our case below). This universal behavior is closely related to that observed in random matrix theory (RMT) [18, 19], which is further argued to be connected to the late-time dynamics of black holes [20].
@@ 2 Krylov Spread Complexity
[…]
@@
[…]
@@
[…]
@@ 3.1 Model Architecture
where $o_{i}$ is the $i$-th component of the output in the final layer L1L-1, which has $N_{(L-1)}$ components. Here $l$ is the layer index, ranging from 1 to L1L-1. $h_{j}^{(l)}$ is the value of node $j$ in layer $l$, which has $N_{(l)}$ nodes. The input layer nodes are represented by $h_{n}^{(0)}=V_{n}$, the $n$-th component of the $N_{(0)}$ input variables. $f_{out}$ is the activation function for the output nodes. This function (for example, $Softmax$, SigmoidSigmoid, or $Linear$) is selected based on the specific problem of interest. Similarly the function $f_{h}$ represents the activation for the nodes $h^{(l)}$ in inner (or hidden) layers 0<l<L10<l<L-1, where usually a nonlinear function such as tanhtanh or Rectified Linear Unit (ReLU) is chosen for the activation. These nonlinear functions are used to introduce nonlinearity into the model, which is essential for learning complex patterns. Specifically, the function ReLU is defined as $f_{\text{ReLU}}(x)=\alpha x$ if $x>0$, otherwise $f_{\text{ReLU}}(x)=0$, with a typical choice of $\alpha>0$. The connection parameters $W$ and the bias $b$, which are randomly initialized in an appropriate manner [24], represent the weights applied to the output vector of the nodes in the previous layer and the corresponding offset. These parameters will be repeatedly updated and optimized during training iteration by minimizing the loss associated with the task at hand.
@@
where $N_{ch}^{(l-1)}$ is the number of channels (kernels) of layer l1l-1, $n_{1}\,(=1,\cdots,X_{(l)})$ and $n_{2}\,(=1,\cdots,Y_{(l)})$ represent the 2D pixel coordinates on the feature maps with sizes of ($X_{(l)},Y_{(l)}$) in layer $l$. In the case of the input layer, $X_{(0)}$ and $Y_{(0)}$ correspond to the number of pixels along $x$ and $y$ axis of the input image. Assuming no padding is used [^1] Padding refers to the process of adding extra pixels, typically zeros, around the borders of an image (or feature map) to control the spatial size reduction caused by convolution operations. In ‘valid’ padding (i.e. no padding), the size of the feature map shrinks, while ‘same’ padding maintains the input size by adding enough pixels to the borders., the feature map size at layer $l$ follows from the relations $X_{(l)}=X_{(l-1)}-K+1$ and $Y_{(l)}=Y_{(l-1)}-K+1$. $m_{1}$ and $m_{2}$ representing convolution indices take integer values ranging from 0 to K1K-1 with the kernel size $K$. Finally, the indices $i$ and $j$ represent the channel index of layer $l$ and l1l-1 respectively.
@@ 3.2 Training
Our model is trained using a supervised learning approach, where the algorithm learns to map inputs to their corresponding outputs based on labeled training data. Specifically, we employ a dataset of N(=1024)N\,(=1024)-dimensional state vectors paired with corresponding quantum complexity values. In this study, the entire dataset consists of $M=200$ samples of GUE, where each sample gives $3N=3072$ state vector data points with $3N$ denoting the total duration of our integer-valued time variable. Among them, 160 samples ( or $160\times 3N$ data points ) are used for training, while $20\times 3N$ data points are used for validation, and another $20\times 3N$ data points are used for testing. The loss function $\mathcal{L}$ used is mean squared error (MSE),
@@
where $C_{I}$ is the predicted quantum complexity value for the $I$-th input data, $\hat{C}_{I}$ is the true quantum complexity value, and batch size $\mathcal{N}$, the number of training examples used in one iteration of the model parameters ($W,b$) update, is set to 32222The batch size 32 (taken as some power of two) is a common choice because it provides a balance between computational efficiency and model convergence. A smaller batch size (like 1) leads to noisy updates and slower convergence, while a larger batch size (equal to the entire dataset) makes the training process inefficient by updating too infrequently. The optimal batch size is usually determined through experimentation, balancing speed and accuracy. . However, when calculating the validation loss, instead of $\mathcal{N}=32$, we used the total number of data points in the validation dataset as the channel number i.e. $\mathcal{N}=20\times 3N=61,440$. Similarly, when calculating the test error, which is the root mean squared error (RMSE), we also use the total number of data points in the test dataset, so $\mathcal{N}=20\times 3N=61,440$. The loss function is minimized with respect to the model parameters $W$ and $b$ using the Adam optimizer [26], an adaptive learning rate optimization algorithm known for its efficiency and effectiveness in training DL models. The model is trained for a total of 100 epochs, where an epoch is one complete pass through the entire training dataset, since the training set has a total of $160\times 3N=491,520$ data points, and the batch size is 32, each epoch corresponds to 15,360 training iterations. In Figure 4, we plot loss functions against epochs for several different bases to examine the convergence behavior of the models during training. The convergence test is an important task in numerical calculations to illustrate the capability of our model and to detect potential problems during training. As can be seen in Figure 4, the stable convergence of loss functions with respect to epochs was obtained for the energy and Krylov bases. Perhaps due to the random nature of the original basis, our training algorithms get stuck immediately. For the pseudo-random basis, both training and validation processes still converge to a much larger value of the loss function compared to those of the energy and Krylov bases. Contrary to expectations, our results suggest that the choice of basis appears to be crucial in applying DL to the complexity.
@@ 4.1 Basis dependence of DL performances
As explained in detail in Section 2.1, the state represented in the original basis for each sample random Hamiltonian becomes completely featureless. Due to the randomness of the basis transformations, our DL algorithms fail to extract the corresponding information on the spread complexity. In the case of the pseudo-random basis, we are simply adding a fixed noise, which is generally known to be effectively handled by DL algorithms. One may also attempt to include each sample Hamiltonian as input data for the DL, but with this inclusion, we find that the aforementioned features of basis dependence do not essentially change.
@@ 5 Conclusions
In this note, we study how a machine based on DL algorithms learns the Krylov spread complexity from patterns of quantum states in systems with $N\times N$ random Hamiltonians drawn from the GUE of matrices. Taking the initial state as the TFD state, we found that the DL algorithm performs well for states at any time, including the plateau region where the states become almost featureless and random. The TFD initial state in our DL setup could be generalized to nonthermal states that span a broader region of the state space.
@@
The deep learning performance of the Krylov operator complexity [12] is of interest, given that its state space structure differs from that of the Krylov spread complexity. Our investigation of the Krylov spread complexity in this note may be extended to other systems, such as random matrix theories with other ensembles, field theories [27, 28, 29], non-chaotic systems [30], and so on. It will be particularly interesting to see whether deep learning algorithms can effectively distinguish different random matrix ensembles or between chaotic and non-chaotic systems. Here, we use $N=1024$, but studying the learning efficiency and the associated scaling behavior as $N$ changes may also be of interest. However, this would involve computational challenges and warrants further investigation.
@@ Acknowledgement
D.B. was supported by the 2024 Research Fund of the University of Seoul. This work was also supported by the UBAI computing resources at the University of Seoul.